2= 5e^-0.35h


Your are trying to find H

Answer by Theo(13342)

2 = 5e^-0.35h
divide both sides of the equation by 5 to get:

2/5 = e^(-.35h)

take natural log of both sides to get:

ln(2/5) = ln(e^(-.35h)

since log(b^x) = x*log(b), your equation becomes:

ln(2/5) = -.35h * (ln(e))

since ln(e) = 1, your equation becomes:

ln(2/5) = -.35h

divide both sides by -.35 to get:

ln(2/5)/(-.35) = h

since ln(2/5) = (-.916290732), your equation becomes:

(-.916290732)/(-.35) = h

simplify to get:

h = 2.61797352

to prove this is correct, replace h in the original equation to get:


2/5 = e^(-.35h) becomes:


2/5 = e^(-.35*2.61797352)

this becomes:

2/5 = e^(-.916290732)

this becomes:

.4 = .4 confirming the answer is correct.